OpenStudy (anonymous):

Convert 46 to base twelve

OpenStudy (anonymous):

0 3 10(A) ------- ------ ----- 12^2 12^1 12^0 3A

OpenStudy (anonymous):

Long divide 12 into 46, and keep the remainder. What do you get?

OpenStudy (anonymous):

How did you get the 3?

OpenStudy (anonymous):

12*3=36 1*10 =10 36+10=46

OpenStudy (anonymous):

In other words, the highest multiple of 12 that \(divides\) evenly into 46 is 3\[\]

OpenStudy (anonymous):

what I meant was , we don't need to do long division

OpenStudy (anonymous):

Fair enough!